# Paying and Persuading

### v4

@@ 7 Discussion
We first show $\mathbb{P}_{\xi}^{\sigma_{n}}\to\mathbb{P}_{\xi}^{\sigma}$. Clearly this is true when $t=0$, since the starting measure is trivial. Inductively, suppose $\mathbb{P}_{\xi-1}^{\sigma_{n}}\to\mathbb{P}_{\xi-1}^{\sigma}$ for every history $h$. Then
$\mathbb{P}_{\xi}^{\sigma_{n}}(h)=\mathbb{P}_{\xi}^{\sigma_{n}}(h^{\xi})=\mathbb{P}_{\xi-1}^{\sigma_{n}}(h^{\xi-1})\sigma_{n}(h^{\xi-1})(x)\implies\mathbb{P}_{\xi-1}^{\sigma}(h^{\xi-1})\sigma(h^{\xi-1})(x)=\mathbb{P}_{\xi}^{\sigma}(h^{\xi})$
@@
$\lim\limits_{a\to a^{*}(\mu_{0})}\left|\frac{\mathbb{E}_{\mu}[v(a,\theta)-v(a^{*}(\mu),\theta)]}{a-a^{*}(\mu_{0})}\right|>\lim\limits_{a\to a^{*}(\mu_{0})}k\left|\frac{\mathbb{E}_{\mu}[u(a,\theta)-u(a^{*}(\mu),\theta)]}{a-a^{*}(\mu_{0})}\right|$
@@
By Proposition 7, it is sufficient to show $\mathcal{V}^{*}(\mu_{0},0)>\text{cav}(V_{0})(\mu_{0})$. Suppose not. Then the static optimal solution without transfers is also the solution to the problem with transfers at $\delta=0$. Let $\{\rho,t,u^{\prime},\alpha\}\in\mathcal{F}(0,\mu_{0})$ be one such solution to the problem where $u^{\prime}\equiv t\equiv 0$, so that $\alpha=a^{*}(\mu,\mathbf{0})$ for all $\mu\in\text{supp}(\rho)$. Because Sender does not do first best, there exists a set $\mathcal{C}$ with $\rho(\mathcal{C})>0$ such that $\frac{\partial}{\partial a}\mathbb{E}_{\mu}[v(a^{*}(\mu,\mathbf{0}),\theta)]\neq 0$ for all $\mu\in\mathcal{C}$. Without loss of generality, suppose this term is positive for all $\mu\in\mathcal{C}$. Define the following actions on $\mathcal{C}$:
@@
$\frac{\mathbb{E}_{\mu}[v(a^{*}(\mu,\mathbf{0})+\varepsilon(\mu),\theta)-v(a^{*}(\mu,\mathbf{0}),\theta)]}{\mathbb{E}_{\mu}[u(a^{*}(\mu,\mathbf{0})+\varepsilon(\mu),\theta)-u(a^{*}(\mu,\mathbf{0}),\theta)]}>k$
@@
The continuum is not without its own challenges. While Proposition 6 and 7 lift immediately for any action space (even those violating Assumption 1), Proposition B.1 does not, and hence it is not immediately clear that Theorem 1 need to hold. Intuitively, this is because the entire space of strategies $\Sigma$ need not be compact absent topological restrictions on its codomain [^20] The cardinality of the codomain is too large to guarantee equicontinuity on the space of functions, so we cannot apply functional Arzela-Ascoli. However, for an arbitrarily fine grid approximation of the continuum, a continuity argument implies that primitives are both incentivizable and we can guarantee that transfers are a last resort.
@@
Consider the interaction between a manager and an employee who must choose between two differentiated projects to pursue. The two projects are differentiated horizontally by the employees’ fit for the project ($\theta\in\{\theta_{0},\theta_{1}\}$) [^21] For example, how much the employee would enjoy being on that project, how complementary their skill sets are to the objective, the difficulty of the project, etc., which the manager privately observes. The manager has a (known) preferred project $a_{1}$ that they would always like the employee to work on, regardless of the characteristics of the other project $a_{0}$ [^22] This can be, for example, because one of the two projects matters for the manager’s promotion package while the other does not.. However, the employee would like to work on whichever project is a better fit for them (e.g. $a_{0}$ at $\theta_{0}$ and $a_{1}$ at $\theta_{1}$), though ex-ante they cannot observe the project’s fit (for example, because the manager can conceal information about each project before they decide which one to join). The employee also always has an outside option, $a_{2}$ (e.g. transferring to a different project), which guarantees some nonzero safe payoff, which is the worst option for the manager. An example of payoffs capturing the above scenario are given below. Fix $k=1$ throughout.
@@
$V(\mu)=\begin{cases}0\text{ for }\mu\in[0,\frac{1}{3})\\ -\frac{1}{2}\text{ for }\mu\in[\frac{1}{3},\frac{2}{3})\\ \frac{5}{2}\text{ for }\mu\in[\frac{2}{3},1]\end{cases}$
@@
Define an operator $\mathcal{O}:\mathcal{B}([-C,C]\times\Delta(\Theta))\to\mathcal{B}([-C,C]\times\Delta(\Theta))$, where $\mathcal{B}(X)$ is the set of bounded functionals from $X\to{R}$. Let $\mathcal{O}$ be defined by
$\mathcal{O}(V)(\bar{u},\mu_{0},\delta)=\max_{\{\rho,t,u^{\prime},\alpha\}}\Big[{E}_{\rho}\Big[(1-\delta)\,{E}_{\mu,\alpha}$ $[v(a,\theta)-kt(\mu,a)]+\delta V(u^{\prime}(\mu,a),M\mu,\delta)\Big]\Big]$
$\text{s.t.}\quad{E}_{\mu,\alpha}[(1-\delta)[u(a,\theta)+t(\mu,a)]+\delta u^{\prime}(\mu,a)]$ $\geq{E}_{\mu}[(1-\delta)u(a^{\prime},\theta)+\delta\underline{U}(\mu)]$
$\hskip 30.00005pt\forall\mu\in\text{supp}(\rho),\;a\in\text{supp}(\alpha(\mu)),\;a^{\prime}\in A$
${E}_{\rho}[(1-\delta)\,{E}_{\mu,\alpha}[u(a,\theta)+t(\mu,a)]+\delta u^{\prime}(\mu,a)]$ $\geq\bar{u}$
${E}_{\rho}[\mu]$ $=\mu_{0}$
$\max\{(1-\delta)t(\mu,a),\,u^{\prime}(\mu,a)\}$ $\in[-C,C]\quad\forall a,\mu$
@@
That $V$ is nonincreasing in $\bar{u}$ follows immediately from the fact that the set of feasible solutions is weakly decreasing in $\bar{u}$ (it can only strictly tighten the promise keeping constraint). Concavity follows by noting that $\mathcal{O}(V)(u,\mu_{0},\delta)$ is concave in $u$ whenever $V$ is concave since the convex combination of a feasible solution at $u$ and $u^{\prime}$ is feasible at $\alpha u+(1-\alpha)u^{\prime}$ (as the constraints are linear in the relevant arguments). Hence, $\mathcal{O}$ maps the set of concave functions to itself, and because the space of concave functions is a closed subset (in the supremum norm) of all functions, we can again apply Banach’s fixed point theorem as before.
@@
Appendix F: Proof of Proposition 8
Proof.
$t^{I}(\mu,a)={E}_{\mu}[u(a^{*}(\mu,\mathbf{0}),\theta)-u(a,\theta)]\mathbf{1}\left\{a\in a^{S}(\mu)\right\}$
where
$a^{S}(\mu)=\operatorname*{\arg\!\max}_{a\in A}{E}_{\mu}[v(a,\theta)-k(u(a^{*}(\mu,\mathbf{0}),\theta)-u(a,\theta))].$
Lemma F.1.
Proof.
${E}_{\mu}[u(a^{*}(\mu,t^{*}),\theta)]+t^{*}(\mu,a^{*}(\mu,t^{*}),)\geq{E}_{\mu}[u(a,\theta)]+t^{*}(\mu,a)\text{ for all }a\in A$
Taking differences implies
${E}_{\mu}[u(a^{*}(\mu,t^{*}),\theta)-u(a,\theta)]\geq t^{*}(\mu,a^{*}(\mu,t^{*}))-t^{*}(\mu,a)\text{ for all }a\in A$
${E}_{\mu}[u(a^{*}(\mu,t^{*}),\theta)-u(a,\theta)]\geq{E}_{\mu}[u(a^{*}(\mu,t^{*}),\theta)-u(a^{*}(\mu,\mathbf{0})),\theta)]=t^{I}\mu)$
Given that canonical transfers are sufficient, it is clear that
$\max_{t}{E}_{\mu}[v(a^{*}(\mu,t),\theta)-kt(\mu,a))]=\max_{a\in A}{E}_{\mu}[v(a,\theta)+u(a,\theta)-u(a^{*}(\mu,\mathbf{0}),\theta)]$
To do so, we need to prove a quick geometric lemma.
Lemma F.2.
$\mathcal{O}_{a}$ is a convex polytope.
Proof.
$V^{t}(\mu)=\max_{a\in A}\left\{{E}_{\mu}[v(a,\theta)-k(u(\tilde{a},\theta)-u(a,\theta))]\right\},$

### v5

@@ 7 Discussion
We first show $\mathbb{P}_{\xi}^{\sigma_{n}}\to\mathbb{P}_{\xi}^{\sigma}$. Clearly this is true when $\xi=0$, since the starting measure is trivial. Inductively, suppose $\mathbb{P}_{\xi-1}^{\sigma_{n}}\to\mathbb{P}_{\xi-1}^{\sigma}$ for every history $h$. Then
$\mathbb{P}_{\xi}^{\sigma_{n}}(h)=\mathbb{P}_{\xi}^{\sigma_{n}}(h^{\xi})=\mathbb{P}_{\xi-1}^{\sigma_{n}}(h^{\xi-1})\sigma_{n}(h^{\xi-1})(x);$
since the inductive hypothesis implies $\mathbb{P}_{\xi-1}^{\sigma_{n}}(h^{\xi-1})\to\mathbb{P}_{\xi-1}^{\sigma}(h^{\xi-1})$ and $\sigma_{n}\to\sigma$ in the strong topology, one has
$\mathbb{P}_{\xi-1}^{\sigma}(h^{\xi-1})\sigma(h^{\xi-1})(x)=\mathbb{P}_{\xi}^{\sigma}(h^{\xi})$
@@
$\lim\limits_{a\to a^{*}(\mu_{0})}\left|\frac{\mathbb{E}_{\mu}[v(a,\theta)-v(a^{*}(\mu_{0}),\theta)]}{a-a^{*}(\mu_{0})}\right|>\lim\limits_{a\to a^{*}(\mu_{0})}k\left|\frac{\mathbb{E}_{\mu}[u(a,\theta)-u(a^{*}(\mu_{0}),\theta)]}{a-a^{*}(\mu_{0})}\right|$
@@
By Proposition 7, it is sufficient to show $\mathcal{V}^{*}(\mu_{0},0)>\text{cav}(V_{0})(\mu_{0})$. Suppose not. Then the static optimal solution without transfers is also the solution to the problem with transfers at $\delta=0$. Let $\{\rho,t,u^{\prime},\alpha\}\in\mathcal{F}(0,\mu_{0})$ be one such solution to the problem where $u^{\prime}\equiv t\equiv 0$, so that $\alpha(\mu)=a^{*}(\mu,\mathbf{0})$ for all $\mu\in\text{supp}(\rho)$. Because Sender does not do first best, there exists a set $\mathcal{C}$ with $\rho(\mathcal{C})>0$ such that $\frac{\partial}{\partial a}\mathbb{E}_{\mu}[v(a^{*}(\mu,\mathbf{0}),\theta)]\neq 0$ for all $\mu\in\mathcal{C}$. Without loss of generality, suppose this term is positive for all $\mu\in\mathcal{C}$. Define the following actions on $\mathcal{C}$:
@@
$\left|\frac{\mathbb{E}_{\mu}[v(a^{*}(\mu,\mathbf{0})+\varepsilon(\mu),\theta)-v(a^{*}(\mu,\mathbf{0}),\theta)]}{\mathbb{E}_{\mu}[u(a^{*}(\mu,\mathbf{0})+\varepsilon(\mu),\theta)-u(a^{*}(\mu,\mathbf{0}),\theta)]}\right|>k$
@@
The continuum is not without its own challenges. While Proposition 6 and 7 lift immediately for any action space (even those violating Assumption 1), Proposition B.1 does not, and hence it is not immediately clear that Theorem 1 need to hold. Intuitively, this is because the entire space of contracts $\Sigma$ need not be compact absent topological restrictions on its codomain [^24] The cardinality of the codomain is too large to guarantee equicontinuity on the space of functions, so one cannot apply functional Arzela-Ascoli. However, for an arbitrarily fine grid approximation of the continuum, a continuity argument implies that primitives are both incentivizable and I can guarantee that transfers are a last resort.
@@
Consider the interaction between a manager and an employee who must choose between two differentiated projects to pursue. The two projects are differentiated horizontally by the employees’ fit for the project ($\theta\in\{\theta_{0},\theta_{1}\}$) [^25] For example, how much the employee would enjoy being on that project, how complementary their skill sets are to the objective, the difficulty of the project, etc., which the manager privately observes. The manager has a (known) preferred project $a_{1}$ that they would always like the employee to work on, regardless of the characteristics of the other project $a_{0}$ [^26] This can be, for example, because one of the two projects matters for the manager’s promotion package while the other does not.. However, the employee would like to work on whichever project is a better fit for them (e.g. $a_{0}$ at $\theta_{0}$ and $a_{1}$ at $\theta_{1}$), though ex-ante they cannot observe the project’s fit (for example, because the manager can conceal information about each project before they decide which one to join). The employee also always has an outside option, $a_{2}$ (e.g. transferring to a different project), which guarantees some safe payoff, normalized to 0, which is the worst option for the manager. An example of payoffs capturing the above scenario are given below. Fix $k=1$ throughout.
@@
$V(\mu)=\begin{cases}0\text{ for }\mu\in[0,\frac{1}{3}]\\ -\frac{1}{2}\text{ for }\mu\in(\frac{1}{3},\frac{2}{3})\\ \frac{5}{2}\text{ for }\mu\in[\frac{2}{3},1]\end{cases}$
@@
Fix $\delta\in(0,1)$. Define an operator $\mathcal{O}:\mathcal{B}([-C,C]\times\Delta(\Theta))\to\mathcal{B}([-C,C]\times\Delta(\Theta))$, where $\mathcal{B}(X)$ is the set of bounded functionals from $X\to{R}$. Let $\mathcal{O}$ be defined by
$\mathcal{O}(V)(\bar{u},\mu_{0},\delta)=\max_{\{\rho,t,u^{\prime},\alpha\}}\Big\{{E}_{\rho}\Big[{E}_{\mu,\alpha}[(1-\delta)\,[v(a,\theta)$ $-kt(\mu,a)]+\delta V(u^{\prime}(\mu,a),M\mu,\delta)]\Big]\Big\}$ (FE)
$\text{s.t.}\quad{E}_{\mu}[(1-\delta)[u(a,\theta)+t(\mu,a)]+\delta u^{\prime}(\mu,a)]$ $\geq{E}_{\mu}[(1-\delta)u(a^{\prime},\theta)+\delta\underline{U}(M\mu)]$ (IC)
${E}_{\rho}[{E}_{\mu,\alpha}[(1-\delta)[u(a,\theta)+t(\mu,a)]+\delta u^{\prime}(\mu,a)]]$ $\geq\bar{u}$ (PK)
${E}_{\rho}[\mu]$ $=\mu_{0}$ (BP)
$\max\{(1-\delta)t(\mu,a),\,|u^{\prime}(\mu,a)|\}\in[0,C]\quad\forall$ $\mu\in\text{supp}(\rho),a\in\text{supp}(\alpha(\mu)),$ (BD)
@@
That $V$ is nonincreasing in $\bar{u}$ follows immediately from the fact that the set of feasible solutions is weakly decreasing in $\bar{u}$ (it can only strictly tighten the promise keeping constraint). Concavity follows by noting that $\mathcal{O}(V)(u,\mu_{0},\delta)$ is concave in $u$ whenever $V$ is concave since the convex combination of a feasible solution at $u$ and $u^{\prime}$ is feasible at $\alpha u+(1-\alpha)u^{\prime}$ (as the constraints are linear in the relevant arguments). Hence, $\mathcal{O}$ maps the set of concave functions to itself, and because the space of concave functions is a closed subset (in the supremum norm) of all functions, one can again apply Banach’s fixed point theorem as before.